OpenStudy (anonymous):

Describe how to transform the quantity of the fifth root of x to the seventh power, to the third powerinto an expression with a rational exponent. Make sure you respond with complete sentences

OpenStudy (anonymous):

@AllTehMaffs do you know how to do this one?

OpenStudy (anonymous):

Is that exactly how the question is written? ... hmm...

OpenStudy (jdoe0001):

\(\bf \textit{keep in mind that}\qquad \huge \sqrt[m]{a^n}=a^{\frac{n}{m}}\)

OpenStudy (jdoe0001):

\(\huge \sqrt[5]{x^{7^3}}\implies \sqrt[5]{x^{49}}\implies \square ?\)

OpenStudy (jdoe0001):

woops

OpenStudy (anonymous):

OH, I get it. Was reading it wrong. So our original expression is \[ \sqrt[5]{(x^7)^3} \] and \[ (x^{a})^b = x^{ab}\] Use that, and then also what jdoe wrote :)

OpenStudy (jdoe0001):

\(\huge \sqrt[5]{x^{7^3}}\implies \sqrt[5]{x^{343}}\implies \square ?\)

OpenStudy (jdoe0001):

wel.... dunno.. is it \(\huge \sqrt[5]{x^{7^3}}\qquad or\qquad \sqrt[5]{(x^7)^3}\quad ?\)

OpenStudy (anonymous):

well @jdoe0001 .... the equation is in parenthesis and the ^3 is on the outside

OpenStudy (anonymous):

then \[ (x^7)^3 = x^{21} \] slightly more manageable :)

OpenStudy (jdoe0001):

\(\huge{ \sqrt[5]{(x^7)^3}\implies \sqrt[5]{x^{7\cdot 3}}\\ \quad \\ \textit{keep in mind that }\quad a^{\frac{n}{m}} = \sqrt[m]{a^n}}\)

OpenStudy (anonymous):

thanks so much! :) @AllTehMaffs @jdoe0001